The minimum energy needed to eject an electron from a potassium atom is 6.96 × 10−19 J. What is the maximum wavelength of light, in nanometers, that will show a photoelectric effect with potassium?
Answer #1

Given:

E = 6.96*10^-19 J

use:

E = h*c/lambda

6.96*10^-19 J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda

lambda = 2.856*10^-7 m

lambda = 286 nm

Answer: 286 nm